Check whether the function f(x,y)={4x^2y/4x^4+y^2 ,(x,y)/=(0,0)
{0,(x,y)=0 is continuous at(0,0)
1
Expert's answer
2020-04-29T16:41:29-0400

Consider the sequence "(x_k, y_k) = (\frac 1k,\frac 1{k^2})" : "f(x_k, y_k)=\frac{\frac 4{k^4}}{\frac 5{k^4}} =\frac 45 \rightarrow \frac 45 \not =0=f(0,0), k \to \infin."

Consequently, the function "f(x,y)" is not continuous at "(0,0)" .